ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 47, 81 Golovanova, Nina F., Golovanov, Anton A., The Moscow State Textile Academy, Russia Eiconal expansion solution for T-matrix equation and the unitarity condition The solution of the T-matrix equation by the eiconal expansion method includes the eiconal parameter. It is defined from the optical theorem to satisfy the unitarity, when the Green function corresponds to Schredinger equation. The other Green function’s cases will be considered and corresponding unitarity conditions will be formulated. Reference: N. F. Golovanova, A. A. Golovanov, Solution of the T-matrixs equation by the eikonal expansion method, ICM, 1998